OpenStudy (anonymous):

When using the substitution method, what expression for x in the second equation would be substituted in the first equation? 4x – 7y = –9 x + 3y = 12

OpenStudy (anonymous):

4x-7y=-9 +4x +4x 7y=4x-9 7y/7=(4x-9)/7 4÷7=4/7 9÷7= 1 2/7 y=4/7x - 1 2/7 (Slope intersept form)